# Enhancing the De-identification of Personally Identifiable Information in Educational Data

### v1

@@ Abstract
Protecting Personally Identifiable Information (PII), such as names, is a critical requirement in learning technologies to safeguard student and teacher privacy and maintain trust. Accurate PII detection is an essential step toward anonymizing sensitive information while preserving the utility of educational data. Motivated by recent advancements in artificial intelligence, our study investigates the GPT-4o-mini model as a cost-effective and efficient solution for PII detection tasks. We explore both prompting and fine-tuning approaches and compare GPT-4o-mini’s performance against established frameworks, including Microsoft Presidio and Azure AI Language. Our evaluation on two public datasets, CRAPII and TSCC, demonstrates that the fine-tuned GPT-4o-mini model achieves superior performance, with a recall of 0.9589 on CRAPII. Additionally, fine-tuned GPT-4o-mini significantly improves precision scores (a threefold increase) while reducing computational costs to nearly one-tenth of those associated with Azure AI Language. Furthermore, our bias analysis reveals that the fine-tuned GPT-4o-mini model consistently delivers accurate results across diverse cultural backgrounds and genders. The generalizability analysis using the TSCC dataset further highlights its robustness, achieving a recall of 0.9895 with minimal additional training data from TSCC. These results emphasize the potential of fine-tuned GPT-4o-mini as an accurate and cost-effective tool for PII detection in educational data. It offers robust privacy protection while preserving the data’s utility for research and pedagogical analysis. Our code is available on GitHub: https://github.com/AnonJD/PrivacyAI
Index Terms:
@@
Privacy, De-identification, Anonymization, Personally Identifiable Information, Large Language Models, Fine-tuning GPT, Cost-effectiveness, Hidden in Plain Sight
I Introduction
@@
Personally Identifiable Information (PII) includes the information (e.g., names, email addresses, and phone numbers) that can identify an individual. Protecting PII in educational data is paramount as learning technologies, including artificial intelligence-powered systems, become increasingly integral to education across all levels. For instance, online human tutoring platforms collect vast amounts of interaction data, including conversational dialogues between students and tutors [1]. While the analysis of such educational data offers insights for data-informed research and improved pedagogical practices [2, 3, 4], it also introduces significant privacy risks due to the sensitive nature of PII [5]. Safeguarding PII is essential to prevent unauthorized access and misuse [6]. Moreover, robust privacy protections are vital for fostering trust among educational stakeholders—students, educators, and parents—and ensuring the responsible adoption of learning technologies [7]. These efforts align with global regulatory frameworks, such as the General Data Protection Regulation (GDPR) and the Family Educational Rights and Privacy Act (FERPA), which mandate the protection of personal data [8].
@@
Given the vast amounts of data collected by learning systems, automating the anonymization of PII in the education domain has become a critical necessity. Previous work has explored rule-based, statistical-based, and neural network-based approaches to detect PII in education-related tasks, such as essay grading [3, 9]. Though showing effectiveness, given the advent of advanced artificial intelligence (AI) models, Pii detection can be further enhanced in detection accuracy, robustness, and generalizability when applied to diverse and complex educational datasets. Moreover, many previous works [10, 11, 12, 13] focus solely on redacting PII from the original data (e.g., the name John from the original sentence “Thanks, John” was processed by redacting “Thanks, [REDACTED]). While this approach mitigates certain risks, it falls short of ensuring comprehensive privacy protection, especially given that current models still fail to achieve perfect accuracy in PII detection [14]. This raises concerns about potential PII leakage. Thus, these challenges underscore the urgent need for more advanced, scalable, and robust methods to facilitate the anonymization process in educational data, ensuring comprehensive protection of PII and mitigating the risks of unintended exposure.
As proposed by [6], the Hidden in Plain Sight (HIPS) method introduces the protection of PII in datasets by first identifying PII entities and then replacing them with synthetic information that retains contextual characteristics. Unlike traditional redaction methods that use generic tokens such as [REDACTED] to redact PII, the first step of the HIPS method replaces PII with tokens that indicate the type of information, such as <Name>, <Email>, or <Address>. The process of identifying PII entities relies on Named Entity Recognition (NER), a task in natural language processing that aims to locate and classify specific entities within text into predefined categories, such as names, email addresses, and phone numbers [15]. Once NER is performed on the data corpus, the original PII entities can be replaced with synthetic information while retaining their contextual relevance. For example, a sentence such as {John Smith}Name lives at {123 Main Street}Address can be transformed into “Alex Doe lives at 456 Elm Avenue,” preserving the sentence structure while anonymizing sensitive details. According to [16], the HIPS method enhances the protection of PII by minimizing the risk of re-identification through context-aware anonymization, while still maintaining the utility of the dataset for research and analysis. However, the effectiveness of the HIPS method also relies on the accuracy of the initial PII identification step performed through NER. To enhance PII protection in educational datasets using the HIPS method, it is crucial to explore advanced AI models capable of improving the performance of the NER process.
The recent advancements in large language models (LLMs) have introduced significant opportunities to enhance PII detection. Recent studies [17] have explored the use of LLMs, such as the GPT-4 model, to identify PII in education datasets. While these models exhibit high recall scores, indicating their ability to identify a broad range of PII entities, their performance often comes at a high computational and financial cost [18, 19]. Moreover, these studies frequently report relatively low precision scores, which poses challenges for further data analysis. Although high recall is essential for safeguarding PII by ensuring that sensitive information is comprehensively identified, we argue that high precision is also important in the context of the HIPS method. Low precision, characterized by a high rate of false positives, can lead to unnecessary replacement of non-PII entities using the HIPS method. This can disrupt the semantic integrity of the dataset, potentially breaking important information relevant to further analysis of educational data. For example, consider the sentence, “The {Newton} method is used for optimization,” which could be replaced with “The {David} method is used for optimization.” In this case, the term “Newton” was incorrectly identified as PII and replaced. Such disruptions might break the mathematical context and hinder further analysis.
@@
II Related Work
II-A Deidentification of Personally Identifiable Information
@@
A significant aspect of the deployment of learning technologies in the actual learning and teaching environment is ensuring data privacy. To address the concern about data privacy, two commonly used deidentification methods are direct redaction and Hidden-In-Plain-Sight (HIPS) [10, 20]. Both approaches aim to safeguard PII by concealing sensitive data but differ significantly in their implementation and impact on downstream data usability.
Redaction is one of the most widely used methods for PII protection in education-related research [12, 21]. In this approach, sensitive information is replaced with generic placeholders, such as “[REDACTED]”, effectively removing it from the dataset. For example, the sentence: “John Smith contacted the office via john.smith@example.com” can be redacted as “[REDACTED] contacted the office via [REDACTED].” While redaction is effective at concealing sensitive information, maintaining consistently high accuracy in PII detection often comes at a significant cost. For instance, the study by [17] employed prompting with the GPT-4 model, which incurs expenses of \$30.00 per 1M input tokens and \$60.00 per 1M output tokens. Given the vast amounts of data collected from learning systems, it is crucial to strike a balance between cost and performance to ensure practical scalability.
The HIPS method offers a more nuanced approach to protecting PII [6]. Instead of simply redacting sensitive information, HIPS replaces PII entities with synthetic but semantically equivalent placeholders that retain the category of the original information [16]. Compared to the redaction method, HIPS could effectively enhance privacy protection. Even if some PII entities are missed during the initial detection step, their replacement with synthetic counterparts ensures no sensitive information is exposed. For example, In [22], the MIST system is used to identify PII entities, after which the HIPS method is used to replace those identified entities with realistic surrogates across two corpora. Following substitution, expert annotators examine the anonymized corpus to detect leaked PII—the entities that MIST failed to identify. On average, only 26.8% of the leaked PII are detected, and 62.8% of the entities considered leaked by human attackers are actually not leaked PII. In [20], Osborne et al. introduce BRATsynthetic as a novel HIPS replacement strategy that leverages a Markov chain–based approach to dynamically substitute PII entities with realistic surrogates. This method notably reduces the risk of PII leakage due to false negatives in PHI detection. For instance, under a 5% false negative error rate, document-level leakage is decreased from 94.2% (using a traditional HIPS replacement) to 57.7%.
@@
It should be noted that the effectiveness of the HIPS method relies heavily on the accuracy of the Named Entity Recognition (NER) step. Misidentified entities (false positives) could lead to unnecessary replacements, disrupting the dataset’s integrity [5, 23]. Conversely, undetected PII (false negatives) could leave sensitive information exposed. Thus, enhancing precision is important. However, existing work in the AI in the education field that focuses on developing Named Entity Recognition (NER) systems for downstream HIPS replacement still demonstrates low precision scores. For example, a study reported a precision score of 0.24 [24]. Given the advantages of using the HIPS method compared to redaction, our study focuses on the application of the HIPS method and particularly aims to enhance the detection of PII entities.
II-B Large Language Models for Data Privacy
@@
Recent advancements in LLMs, such as GPT-4, which are trained on extensive datasets from diverse domains, enable them to capture long-range dependencies and contextual nuances that are crucial for identifying PII. Studies leveraging GPT-based models have demonstrated their ability to achieve high recall scores, effectively identifying a broad range of PII entities [25]. Additionally, LLMs can leverage this internal knowledge to distinguish actual PII from non-PII. Some recent evidence supports this view: for instance, a recent study prompts GPT-3 for a NER task, suggesting that fine-tuning such large models could yield more notable results [25], especially since fine-tuned LLMs tend to outperform their prompted counterparts on NLP tasks [26, 27].
@@
The approach of fine-tuning an LLM for PII identification aligns well with the emerging AI-for-education domain, where labeled text data for PII identification are often scarce and fragmented. Due to this limited availability, a model can be trained on data that are not representative of the actual use case, and less experienced users may require a tool that can quickly adapt to their specific domain. Consequently, a source model capable of learning effectively from sparse and unrepresentative datasets is needed. Previous work has shown that an LLM can be fine-tuned with just a few labeled examples (few-shot learning) [27]. Although studies also indicate that successful few-shot learning is achievable with other architectures, these typically require carefully designed learning strategies and meticulous parameter tuning to prevent overfitting [28, 29]. In contrast, larger LLMs have been proven to be more resistant to overfitting as their size increases [27, 30], potentially owing to their memorization dynamics [31].
A final advantage of fine-tuning an LLM for PII de-identification lies in its lower financial and computational cost compared to prompt-engineering an LLM for PII de-identification, which often requires multiple demonstrations in the input prompts [25]. The above observation highlights the potential for fine-tuning large models, with GPT emerging as a promising candidate. Several studies have utilized prompted GPT approaches to de-identify PII entities, demonstrating encouraging results [17, 25]. With the growing availability of cost-effective fine-tuning APIs  [^1] https://openai.com/api/pricing/ and the lack of prior work using fine-tuned GPT for PII identification, we aim to explore this approach further.
III Methods
III-A Data and Data Pre-processing
@@
Our study utilized the Cleaned Repository of Annotated Personally Identifiable Information (CRAPII) [^2] Cleaned Repository of Annotated PII. https://www.kaggle.com/datasets/langdonholmes/cleaned-repository-of-annotated-pii/data dataset [14], which comprises 22,688 samples of student writings collected from a massive open online course (MOOC) offered by a university in the United States. The course focused on critical thinking through design, teaching learners strategies such as storytelling and visualization to solve real-world problems [14]. The dataset includes seven PII categories as direct identifiers: Names, Email Addresses, Usernames, IDs, Phone Numbers, Personal URLs, and Street Addresses [14]. A sample of the dataset is shown in Table I. In total, the dataset contains 4,871 labeled words categorized as PII entities. To enable entity-based matching during our analysis, we extracted the character-wise positions of all annotated PII entities within the text.
TABLE I: Illustrated Example of CRAPII Dataset
@@
full_text Hi John Doe. Tel: (555)555-$5$
@@
‘(555)555-$5$’]
labels [‘O’, ‘B-NAME’,
‘I-NAME’, ‘O’, ‘O’,
‘O’, ‘B-PHONE_NUM’]
@@
We also introduced another dataset, the Teacher-Student Chatroom Corpus (TSCC) [10], to examine the generalizability of our investigated models, as detailed in Section III-F. Generalizability is crucial for evaluating a model’s robustness when applied to datasets with different distributions. The TSCC dataset contains 260 chatroom sessions, with a total of 41.4K conversational turns and 362K word tokens. We processed the dataset by extracting the role and edited columns and combining them into a simplified role: text format, where each conversational turn is represented on a new line. A sample excerpt of the processed transcription is shown below:
@@ III-B Models for PII Detection
Microsoft Presidio. [^3] Microsoft Presidio: Data Protection and De-identification SDK. https://microsoft.github.io/presidio/ It is an open-source toolkit designed for detecting and anonymizing PII across various text and image formats. It offers pre-built NER (Named Entity Recognition) models, such as en_core_web_lg and en_core_web_trf, which utilize linguistic patterns and deep learning techniques to classify words or phrase into predefined named entities (e.g., names, emails, and phone numbers). Additionally, Presidio provides options for integrating external machine learning models for enhancing PII detection.
@@
In our study, Presidio serves as one of the baseline models for detecting PII entities. As indicated in previous work [14], en_core_web_lg is a standard NER model based on the spaCy [^4] https://spacy.io/ framework that primarily relies on statistical techniques, while the transformer-based en_core_web_trf leverages pre-trained transformers, capturing long-range dependencies for more accurate context-based entity recognition. Thus, we primarily used these two configurations: en_core_web_lg and en_core_web_trf, as depicted in Fig. 1 (①).
@@
Azure AI Language. [^5] Microsoft Azure AI Language: Personally Identifiable Information (PII) Detection. https://learn.microsoft.com/en-us/azure/ai-services/language-service/personally-identifiable-information/overview It offers a cloud-based PII detection service capable of identifying and redacting sensitive information such as phone numbers and email addresses. Our study adopted Azure AI Language as one of the baseline models for detecting PII entities in our dataset, as illustrated in Fig. 1 (②). We used asynchronous processing through the REST API [^6] Azure AI Language Service: Using Asynchronous Requests. https://learn.microsoft.com/en-us/azure/ai-services/language-service/concepts/use-asynchronously to handle texts up to 125,000 characters per document. This approach is necessary because the longest transcript in our dataset contains 17,405 characters, exceeding the synchronous processing limit of 5,120 characters. Requests are batched with 5 documents per request and a rate limit of 1000 requests per minute, ensuring compliance with Azure’s service limits for PII detection [^7] Azure AI Language Service: Data Limits. https://learn.microsoft.com/en-us/azure/ai-services/language-service/concepts/data-limits. Each asynchronous request was completed within 24 hours, ensuring timely processing of all data.
@@
GPT-4o-mini (Prompting). Motivated by the effectiveness of prompting LLMs in de-identifying PII, as demonstrated in a recent study [17], our study adapts their prompting strategy with modifications to fit our task. Instead of employing GPT-4, as used in their work [17], we opted for GPT-4o-mini, which requires only 1/200 of the cost per million input tokens and 1/100 of the cost per million output tokens compared to GPT-4888OpenAI Models Pricing: https://openai.com/api/pricing/. Then, we modified the prompt structure by leveraging special identifiers for labeling detected entities rather than using a redaction method such as replacing detected PII with a generic label like [REDACTED] as shown in their work [17].
@@
Inspired by the GPT-NER method [25], special identifiers can be used to mark entities, preserving non-PII content and ensuring precise PII detection. This method also reduces issues such as hallucinations and over-labeling that can arise with generative models. In particular, we require the detected entity positions to be an exact match to the true entity positions, but GPT often struggles with positional accuracy in long texts due to hallucinations and counting limitations [32]. To address this, we let GPT label detected PII entities by surrounding them with special identifiers for different categories, as outlined in Table II, and subsequently extract them using regular expressions, ensuring accurate detection and positioning.
TABLE II: Special Tokens for PII Detection
@@
Phone Number %%Text ~~~ %%%(555)555-$5$~~~
We employ a few-shot learning strategy to guide GPT-4o-mini in accurately identifying different types of PII entities. Few-shot learning provides contextual examples that help the model understand the diversity of entity types and nuances within our dataset [33]. This approach is chosen because the additional examples often improve the model’s ability to handle ambiguous cases and enhance consistency in PII identification. Table III outlines the structure of the System, User, and Assistant prompts used in prompting GPT-4o-mini. Three examples [^9] https://github.com/AnonJD/PrivacyAI/blob/main/Privacy_main/prompts.txt are selected from the CRAPII dataset and incorporated into the user prompt. The Assistant’s output is the input text with labeled PII entities. The prompting process is detailed in Fig. 1 (③).
TABLE III: Prompt Structures for Prompting-GPT-4o-mini Using Special Identifiers
@@
System You are an expert in labeling Personally Identifiable Information (PII). Start your response right away without adding any prefix (such as “Response:”) or suffix. Use special identifiers to mark different types of PII in the given text.
@@
GPT-4o-mini (Fine-tuning). Recent studies [4] have demonstrated the effectiveness of fine-tuning LLMs for NER tasks, specifically showing that fine-tuned LLMs significantly outperform prompting-based approaches in accurately identifying entities. Inspired by these findings, our study employed a fine-tuning strategy on GPT-4o-mini for PII detection. We utilized the same approach of incorporating special tokens, as shown in Table II. Then, GPT-4o-mini was fine-tuned on the training dataset and evaluated on the test dataset to assess its performance. The detailed structure of the system, user, and assistant prompts used during the fine-tuning process is presented in Table IV, while the implementation process is illustrated in Fig. 1 (④).
TABLE IV: Details of Fine-tuning GPT-4o-mini Using Special Identifiers
@@
TABLE V: Prompt Structures for Verifier Models

### v2

@@ Abstract
Protecting Personally Identifiable Information (PII), such as names, is a critical requirement in educational data mining to safeguard the privacy of students and teachers and maintain trust. Accurate PII detection is an essential step toward anonymizing sensitive information while preserving the utility of educational data. Motivated by recent advancements in artificial intelligence, our study investigates the GPT-4o-mini model as a cost-effective and efficient solution for PII detection tasks. We explore both prompting and fine-tuning approaches and compare GPT-4o-mini’s performance against established frameworks, including Microsoft Presidio and Azure AI Language. Our evaluation on two public datasets, CRAPII and TSCC, demonstrates that the fine-tuned GPT-4o-mini model achieves superior performance, with a recall of 0.9589 on CRAPII. Additionally, fine-tuned GPT-4o-mini significantly improves precision scores (a threefold increase) while reducing computational costs to nearly one-tenth of those associated with Azure AI Language. Furthermore, our bias analysis reveals that the fine-tuned GPT-4o-mini model consistently delivers accurate results across diverse cultural backgrounds and genders. The generalizability analysis using the TSCC dataset further highlights its robustness, achieving a recall of 0.9895 with minimal additional training data from TSCC. These results emphasize the potential of fine-tuned GPT-4o-mini as an accurate and cost-effective tool for PII detection in educational data. It offers robust privacy protection while preserving the data’s utility for research and pedagogical analysis. Our code is available on GitHub: https://github.com/AnonJD/PrivacyAI
@@
Keywords: privacy, de-identification, personally identifiable information, large language models, fine-tuning GPT, cost-effectiveness, hidden in plain sight
1 Introduction
@@
Personally Identifiable Information (PII) includes the information (e.g., names, email addresses, and phone numbers) that can identify an individual. Protecting PII is critical as computer-assisted learning systems become increasingly pervasive across all educational levels. For example, online human tutoring platforms collect vast amounts of interaction data, including conversational dialogues between students and tutors [Reidenberg and Schaub (2018)]. While the analysis of such educational data offers insights for data-informed research and improved pedagogical practices [Portnoff et al. (2021), Chen et al. (2016), Lin et al. (2024)], it also introduces significant privacy risks due to the sensitive nature of PII [Fung et al. (2010)]. Safeguarding PII is essential to prevent unauthorized access and misuse [Carrell et al. (2012)]. Moreover, robust privacy protections are vital for fostering trust among educational stakeholders—students, educators, and parents—and ensuring the responsible adoption of learning technologies [Zeide (2018)]. These efforts align with global regulatory frameworks, such as the General Data Protection Regulation (GDPR) and the Family Educational Rights and Privacy Act (FERPA), which mandate the protection of personal data.
@@
Given the vast amounts of data collected by learning systems, automating the anonymization of PII in the education domain has become a critical necessity. Previous work has explored rule-based, statistical-based, and neural network-based approaches to detect PII in education-related tasks, such as essay grading [Chen et al. (2016), Liu et al. (2023)]. Though showing effectiveness, given the advent of advanced AI models, PII detection can be further enhanced in detection accuracy, robustness, and generalizability when applied to diverse and complex educational datasets. Moreover, many previous works [Caines et al. (2022), Pal et al. (2024), Lison et al. (2021), Samarati (2001)] focus solely on redacting PII from the original data (e.g., the name John from the original sentence “Thanks, John” processed by redacting: “Thanks, [REDACTED]). While this approach mitigates certain risks, it falls short of ensuring comprehensive privacy protection, especially given that current models still fail to achieve perfect accuracy in PII detection [Holmes et al. (2024)]. This raises concerns about potential PII leakage. Thus, these challenges underscore the urgent need for more advanced, scalable, and robust methods to facilitate the anonymization process in educational data, ensuring comprehensive protection of PII and mitigating the risks of unintended exposure.
[…]
@@
By answering this question, our study makes the following key contributions:
(1) Fine-Tuned Model for PII Detection: We employ and evaluate a fine-tuned GPT-4o-mini model for PII detection. Our findings show it achieves high recall scores (over 0.95) at a fraction of the cost of comparable commercial services.
(2) Verifier Model for Enhancing Precision: We introduce a two-step verifier model that enhances detection precision. By reducing false positives, this approach helps better preserve non-PII data, which can support de-identification methods such as Hidden in Plain Sight (HIPS) that aim to minimize disruption to non-sensitive information.
(3) Cultural and Gender Bias Analysis: We conduct an analysis of model fairness, evaluating performance across cultural and gender groups using name-based subgroup analysis in the TSCC dataset, where names were replaced according to cultural and gender distributions. Our results reveal that the fine-tuned GPT-4o-mini delivers equitable performance, reducing the cultural biases present in established baseline models.
(4) Evaluation on Domain Adaptation Study: We demonstrate the model’s robustness by testing it on a different dataset (i.e., TSCC). The model proves to be adaptable, achieving high accuracy (Precision of 0.9708, Recall of 0.9895, and $F_{1}$ Score of 0.9801) on the new domain after fine-tuning on a small sample, confirming its suitability for deployment across varied educational contexts.
2 Related Work
2.1 Deidentification of Personally Identifiable Information
@@
A significant aspect of the deployment of learning technologies in the actual learning and teaching environment is ensuring data privacy. To address the concern about data privacy, two commonly used deidentification methods are direct redaction and Hidden In Plain Sight (HIPS) [Caines et al. (2022), Osborne et al. (2022)]. Both approaches aim to safeguard PII by concealing sensitive data but they differ significantly in their implementation and impact on downstream data usability.
[…]
@@
It should be noted that the effectiveness of the HIPS method relies heavily on the accuracy of the Named Entity Recognition (NER) step. Misidentified entities (false positives) could lead to unnecessary replacements, disrupting the dataset’s integrity [Fung et al. (2010), Carvalho et al. (2022)]. Conversely, undetected PII (false negatives) could leave sensitive information exposed. Thus, enhancing precision is important. However, existing work in the AI in education field that focuses on developing NER systems for downstream HIPS replacement still demonstrates low precision scores. For example, one study reported a precision score of 0.24 [Holmes et al. (2023)]. Given the advantages of using the HIPS method compared to redaction, our study focuses on the application of the HIPS method and particularly aims to enhance the detection of PII entities.
2.2 Large Language Models for Data Privacy
@@
Recent advancements in LLMs, such as GPT-4, which are trained on extensive datasets from diverse domains, enable them to capture long-range dependencies and contextual nuances that are crucial for identifying PII. Studies leveraging GPT-based models have demonstrated their ability to achieve high recall scores, effectively identifying a broad range of PII entities [Wang et al. (2025)]. Additionally, LLMs can leverage this internal knowledge to distinguish actual PII from non-PII. Some recent evidence supports this view: for instance, a recent study prompts GPT-3 for a NER task, suggesting that fine-tuning such large models could yield more notable results [Wang et al. (2025)], especially since fine-tuned LLMs tend to outperform their prompted counterparts on NLP tasks [Zhang et al. (2023), Mosbach et al. (2023)].
@@
The approach of fine-tuning an LLM for PII identification aligns well with the emerging AI-for-education domain, where labeled text data for PII identification are often scarce and fragmented. Due to this limited availability, a model can be trained on data that are not representative of the actual use case, and less experienced users may require a tool that can quickly adapt to their specific domain. Consequently, a source model capable of learning effectively from sparse and unrepresentative datasets is needed. Previous work has shown that an LLM can be fine-tuned with just a few labeled examples (few-shot learning) [Mosbach et al. (2023)]. Although studies also indicate that successful few-shot learning is achievable with other architectures, these typically require carefully designed learning strategies and meticulous parameter tuning to prevent overfitting [Shen et al. (2021), Lee et al. (2023)]. In contrast, larger LLMs have been proven to be more resistant to overfitting as their size increases [Mosbach et al. (2023), Gadre et al. (2025)], potentially owing to their memorization dynamics [Tirumala et al. (2022)].
There has been work showing that current standard redactors perform worse in certain gender and cultural groups. The redactors identify names in the African and Asian / Pacific cultural group with a higher error rate [Mansfield et al. (2022)]. Performance disparity has also been reported between identifying male and female names of an NER system [Zhao et al. (2024)]. LLMs, which have been aligned to remedy gender and cultural bias, could be the base models for the development of a PII identification system with reduced bias [Ouyang et al. (2022)].
[…]
3 Methods
3.1 Data and Data Pre-processing
@@
Our study utilized the Cleaned Repository of Annotated Personally Identifiable Information (CRAPII) [^1] Cleaned Repository of Annotated PII. https://www.kaggle.com/datasets/langdonholmes/cleaned-repository-of-annotated-pii/data dataset [Holmes et al. (2024)], which comprises 22,688 samples of student writings collected from a massive open online course (MOOC) offered by a university in the United States. The course focused on critical thinking through design, teaching learners strategies such as storytelling and visualization to solve real-world problems [Holmes et al. (2024)]. The dataset includes seven PII categories as direct identifiers: Names, Email Addresses, Usernames, IDs, Phone Numbers, Personal URLs, and Street Addresses [Holmes et al. (2024)]. A sample of the dataset is shown in Table 1. In total, the dataset contains 4,871 labeled words categorized as PII entities. To enable entity-based matching during our analysis, we extracted the character-wise positions of all annotated PII entities within the text.
Table 1: Illustrated Example of CRAPII Dataset
@@
full_text Hi John Doe. Tel: (555)555-5555
@@
‘(555)555-5555’]
labels [‘O’, ‘B-NAME’, ‘I-NAME’, ‘O’, ‘O’, ‘O’, ‘B-PHONE_NUM’]
@@
We also introduced another dataset, the Teacher-Student Chatroom Corpus (TSCC) [Caines et al. (2022)], to examine the generalizability of our investigated models, as detailed in Section 3.6. Generalizability is crucial for evaluating a model’s robustness when applied to datasets with different distributions. The TSCC dataset contains 260 chatroom sessions, with a total of 41.4K conversational turns and 362K word tokens. We processed the dataset by extracting the role and edited columns and combining them into a simplified role: text format, where each conversational turn is represented on a new line. A sample excerpt of the processed transcription is shown below:
@@ 3.2 Models for PII Detection
[…]
Diagram showing five PII detection models: Presidio, Azure AI Language, and Prompting LLM with arrows indicating workflows.
@@
Microsoft Presidio. [^2] Microsoft Presidio: Data Protection and De-identification SDK. https://microsoft.github.io/presidio/ It is an open-source toolkit designed for detecting and anonymizing PII across various text and image formats. It employs a regular-expression parser alongside NER models (such as en_core_web_lg and en_core_web_trf). Those NER models utilize linguistic patterns and deep learning techniques to classify words or phrases into predefined named entities (e.g., names, emails, and phone numbers). The backbone for en_core_web_lg is a convolutional neural network (CNN), and en_core_web_trf is a transformer-based model. [^3] https://spacy.io/models The model sizes for both are 382 and 436 MB. Additionally, Presidio provides options for integrating external machine learning models for enhancing PII detection.
@@
In our study, Presidio serves as one of the baseline models for detecting PII entities. As indicated in previous work [Holmes et al. (2024)], en_core_web_lg is a standard NER model based on the spaCy [^4] https://spacy.io/ framework that primarily relies on statistical techniques, while the transformer-based en_core_web_trf leverages pre-trained transformers, capturing long-range dependencies for more accurate context-based entity recognition. Thus, we primarily used these two configurations: en_core_web_lg and en_core_web_trf, as depicted in Fig. 1 (①).
@@
Azure AI Language. [^5] Microsoft Azure AI Language: Personally Identifiable Information (PII) Detection. https://learn.microsoft.com/en-us/azure/ai-services/language-service/personally-identifiable-information/overview It offers a cloud-based PII detection service capable of identifying and redacting sensitive information such as phone numbers and email addresses. Our study adopted Azure AI Language as one of the baseline models for detecting PII entities in our dataset, as illustrated in Fig. 1 (②). The version we use is 2023-04-15-preview. We used asynchronous processing through the REST API [^6] Azure AI Language Service: Using Asynchronous Requests. https://learn.microsoft.com/en-us/azure/ai-services/language-service/concepts/use-asynchronously to handle texts up to 125,000 characters per document. This approach is necessary because the longest transcript in our dataset contains 17,405 characters, exceeding the synchronous processing limit of 5,120 characters. Requests are batched with 5 documents per request and a rate limit of 1000 requests per minute, ensuring compliance with Azure’s service limits for PII detection. [^7] Azure AI Language Service: Data Limits. https://learn.microsoft.com/en-us/azure/ai-services/language-service/concepts/data-limits Each asynchronous request was completed within 24 hours, ensuring timely processing of all data.
@@
GPT-4o-mini (Prompting). Motivated by the effectiveness of prompting LLMs in de-identifying PII, as demonstrated in a recent study [Singhal et al. (2024)], our study adapts their prompting strategy with modifications to fit our task. Instead of employing GPT-4, as used in their work [Singhal et al. (2024)], we opted for GPT-4o-mini, which requires only 1/200 of the cost per million input tokens and 1/100 of the cost per million output tokens compared to GPT-4.888OpenAI Models Pricing: https://openai.com/api/pricing/ Then, we modified the prompt structure by leveraging special identifiers for labeling detected entities rather than using a redaction method such as replacing detected PII with a generic label like [REDACTED] as shown in their work [Singhal et al. (2024)].
@@
Inspired by the GPT-NER method [Wang et al. (2025)], special identifiers can be used to mark entities, preserving non-PII content and ensuring precise PII detection. This method also reduces issues such as hallucinations and over-labeling that can arise with generative models. In particular, we require the detected entity positions to be an exact match to the true entity positions, but GPT often struggles with positional accuracy in long texts due to hallucinations and counting limitations [Ball et al. (2024)]. To address this, we let GPT label detected PII entities by surrounding them with special identifiers for different categories (see Table 2) and subsequently extract them using regular expressions, ensuring accurate detection and positioning.
Table 2: Special Tokens for PII Detection
@@
Phone Number %%%Text~~~ %%%(555)555-5555~~~
[…]
Table 3: Prompt Structures for Prompting-GPT-4o-mini Using Special Identifiers
@@
System You are an expert in labeling Personally Identifiable Information (PII). Start your response right away without adding any prefix (such as “Response:”) or suffix.
@@
GPT-4o-mini (Fine-tuning). Our study employed a fine-tuning strategy on GPT-4o-mini for PII detection. We utilized the same approach of incorporating special tokens, as shown in Table 2. Then, GPT-4o-mini was fine-tuned on the Base Train Set and evaluated on the Test Set (as defined in Section 3.3) to assess its performance. The detailed structure of the system, user, and assistant prompts used during the fine-tuning process is presented in Table 4, while the implementation process is illustrated in Fig. 1 (④).
Table 4: Details of Fine-tuning GPT-4o-mini Using Special Identifiers
@@
Verifier Models. To improve the precision (Equation 1) of PII detection while maintaining recall (Equation 2), we propose the addition of a verifier model as a second step to verify whether predictions are accurately identified as PII. This approach is inspired by recent studies [Brandfonbrener et al. (2024), Li et al. (2023)] that integrate verifiers into multi-step reasoning tasks in LLMs. For the implementation of verifier models, we propose two variants: Verifier Model I (Without CoT) and Verifier Model II (With CoT). These verifier models assess detected entities within their surrounding context to eliminate false positives while retaining true PII entities, thereby enhancing the precision of PII detection systems. Both verifiers are fine-tuned versions of the GPT-4o-mini model. The process of the verifier model approach is illustrated in Fig. 1 (⑤).
The Verifier Model II (With CoT) incorporates Chain-of-Thought (CoT) reasoning, a method shown to enhance decision-making by breaking down complex problems into intermediate steps [Wei et al. (2022)]. The Verifier Model II (With CoT) aims to improve interpretability and robustness by generating reasoning before the final classification. However, it generates more output tokens, leading to higher computational costs. The Verifier Model I (Without CoT), which directly classifies entities without reasoning, is retained for scenarios where computational resources are limited or speed is a higher priority.
To construct the training data for the verifier models, we used the dedicated Verifier Train Set (as defined in Section 3.3). The process begins by running a base PII detection model of the user’s choice on this dataset to generate an initial list of detected PII entities. For each detected entity from this list, we then extracted it along with its surrounding textual context. This (entity, context) pair was subsequently labeled as T (True PII) or F (False PII) by comparing the entity to the ground-truth labels.
For the Verifier Model II (With CoT), we use GPT-4o-mini to generate reasoning to support the classification. If the reasoning does not align with the ground truth after six attempts, the label defaults to T to avoid mistakenly removing true PII entities and prioritize privacy preservation. The prompt structures for both verifier models are shown in Table 5.
Table 5: Prompt Structures for Verifier Models